### v1

# Evaluating World Models with LLM for Decision Making

Abstract
World model emerges as a key module in decision making, where MuZero and Dreamer achieve remarkable successes in complex tasks. Recent work leverages Large Language Models (LLMs) as general world simulators to simulate the dynamics of the world due to their generalizability. LLMs also serve as the world model for deliberative reasoning in Reasoning via Planning (RAP) and Tree of Thought (ToT). However, the world models are either evaluated as a general world simulator, or as a functional module of the agent, i.e., predicting the transitions to assist the planning. In this work, we propose a comprehensive evaluation of the world models with LLMs from the decision making perspective. Specifically, we leverage the 31 diverse environments from (Wang et al., 2023; 2024) and curate the rule-based policy of each environment for the diverse evaluation. Then, we design three main tasks, i.e., policy verification, action proposal, and policy planning, where the world models can be used for decision making solely. Finally, we conduct the comprehensive evaluation of the advanced LLMs, i.e., GPT-4o and GPT-4o-mini, on the environments for the three main tasks under various settings. The key observations include: i) GPT-4o significantly outperforms GPT-4o-mini on the three main tasks, especially for the tasks which require the domain knowledge, ii) the performance of the world model with LLM will be decreased for long-term decision-making tasks, and iii) the combination of different functionalities of the world model will brings additional unstabilities of the performance.
1 Introduction
Due to the celebrating success of MuZero (Schrittwieser et al., 2020) and Dreamer (Hafner et al., 2019; 2021; 2023), world model (Ha & Schmidhuber, 2018) becomes a key concept in decision making. With the encoding vast knowledge of the world for the prediction of the new states after taking the actions, world models have demonstrate their effectiveness for generalizing to novel tasks (Byravan et al., 2020), efficient planning with world model (Sekar et al., 2020), and working on offline datasets (Schrittwieser et al., 2021; Yu et al., 2020; 2021). World models can also be viewed as general world simulators where users can interact with, e.g., Genie (Bruce et al., 2024) and Vista (Gao et al., 2024).
Large Language Models (LLMs) achieve remarkable success in enormous natural language tasks in the past five years (Brown et al., 2020; OpenAI, 2023). Several recent works leverage LLMs as the generalist world models to provide the environment knowledge for various complex tasks, e.g., math and reasoning. With the fine-tuning over pre-collected data from the environments, the LLMs can predict the action sequences across different tasks over environments while maintaining the capabilities on other domains (Xiang et al., 2023). LLMs also serve as the world model explicitly in Reasoning via Planning (RAP) (Hao et al., 2023) and Reason for Future, Act for Now (RAFA) (Liu et al., 2023), where the LLMs predict the next states based on the actions executed at current states, e.g., the states of blocks in the BlocksWorld (Valmeekam et al., 2023), which is used to assist the planning methods. On the other hand, LLMs serve as the world model implicitly in the widely-used Tree of Thoughts (ToT) (Yao et al., 2023), as well as Graph of Thoughts (GoT) (Besta et al., 2024), where the LLMs need to predict the states and evaluate the thoughts. Recent work also consider LLMs as world simulators (Wang et al., 2024; Xie et al., 2024), where they evaluate the performance of LLMs on the prediction of next states and the game progress, demonstrating the potentials of LLMs as general world models.
Figure 1: Picking the action with higher value. The prediction of B in Case 2 is more accurate than Case 1 in term of the $L_{2}$ loss, but leads to the wrong action.
However, most of the previous works evaluate the world models with LLMs either as general world simulators (Wang et al., 2024), or as additional modules of the agents to make decisions (Liu et al., 2023). We argue that a comprehensive evaluation of the world models themselves from the decision-making perspective is needed due to the two facts. First, the objective of decision making is finding the policy to complete the task, where only a small portion of the world will be visited during finding and executing the policy, given the fact that AlphaZero can find the super-human policy (Silver et al., 2018) by only exploring a small proportion (less than 1%) of the state space, therefore, the evaluation of the world models on the predictions of the transitions that relevant to the desired policy is more important than the transitions far from the policy. Second, the influence of the world models is usually coupled with the actors who choose the actions, i.e., if the actor cannot pick the correct actions, the task cannot be completed even when the world model is accurate, which brings additional difficulties to understand the performances of the world models.
Therefore, we address these issues with three key observations in this work:
  • Observation 1: Prediction is important, but not that important. An illustrative example is displayed in Figure 1, which indicates that more accurate predictions do not lead to right decisions. [^1] The issue in Figure 1 can be elicited by various methods, e.g., rank prediction. This example is just to illustrate the discrepancy between prediction and decision, motivating us to reconsider the evaluation of the world model for decision making. This phenomenon is also observed in other decision making scenarios, e.g., financial trading (Sun et al., 2023). The success of MuZero Unplugged (Schrittwieser et al., 2021) also demonstrate that we can learn good policies from inaccurate world models which are trained only with limited data. This motivates us that the evaluation of the world models for decision making should focus on the predictions which relevant to the desired policy, rather than as general world simulators.
  • Observation 2: Selecting potential actions should be an important feature for world models. Most of the previous works in world model focus on next state and reward prediction, and the action selection is usually completed by separate model. Several works (Xie et al., 2024; Janner et al., 2021) incorporate the action selection into consideration. We argue that as the world model for decision making has more knowledge about the world, which can potentially make a better selection of the potential actions. World models can also be viewed as game engines (Valevski et al., 2024), which have to provide potential actions to guide fresh players to complete tasks, e.g., Red Dead Redemption 2 (Tan et al., 2024).
  • Observation 3: Planning with world models can find the policies solely. With the prediction of the next states and the action proposal, we can leverage planning methods or search methods to find the policies. Most works introduce the critic (i.e., the value function) to evaluate the actions immediately for efficient planning (Schrittwieser et al., 2020; Hao et al., 2023), however, the critic is not necessary for finding policies and may also influence the performance. Therefore, we should evaluate the world model solely for finding the policies to avoid the influences of other modules.
Figure 2: Evaluation of World Model with LLM for Decision Making.
Therefore, based on the above three key observations, we proposal a comprehensive evaluation of world models with LLMs for decision making. Specifically, we leverage 31 diverse environments from (Wang et al., 2023; 2024) with different tasks varying from daily tasks, e.g., washing clothes, to scientific tasks, e.g., forging keys, and different difficulty levels, i.e., steps to complete the tasks, and curate the rule-based policy for each environment for the evaluation. Then, we design three main tasks: i) policy verification: verifying whether the policy can complete the task, ii) action proposal: proposing the top-$K$ actions that can potentially complete the task, and iii) policy planning: finding the policy solely with the combination of the different functionalities, i.e., policy verification and action proposal. Finally, we conduct the comprehensive evaluation of the advanced LLMs, i.e., GPT-4o and GPT-4o-mini, on the environments for the three tasks under various settings. The key observations include: i) GPT-4o significantly outperforms GPT-4o-mini on the three main tasks, especially for the tasks which requires the domain knowledge, ii) the performance of the world model with LLM will be decreased for long-term decision-making tasks, and iii) the combination of different functionalities of the world model for decision making will brings unstability of the performance.
2 Related Work
In this work, we provide a detail review of the related work, which is divided into three main categories, i.e., world models in decision making, LLMs as world simulators, world models in LLMs.
World Models in Decision Making.
World models are actively explored by researchers to further improve the agent’s performance and the sample efficiency (Ha & Schmidhuber, 2018; Janner et al., 2019; Hafner et al., 2019; Schrittwieser et al., 2020). Dreamer (Hafner et al., 2019) is a practical model-based reinforcement learning algorithm that introduces the belief over states as a part of the input to the model-free DRL algorithm used. MuZero (Schrittwieser et al., 2020) is a remarkable success of model-based RL, which learns the world model and conduct the planning in the latent space. MuZero achieves the superior performances over other model-based and model-free RL methods. The world models trained in these methods are problem-specific and cannot be generalized to other problems, which motivates researchers to seek to more generalizable world models, e.g., LLMs as world models. The world model with LLM in (Xiang et al., 2023) is trained to gain the environment knowledge, while maintaining other capabilities of the LLMs. Dynalang (Lin et al., 2024) proposes the multimodal world model, which unifies the videos and texts for the future prediction in decision making.
LLMs as World Simulators.
World simulators are developed to model the dynamics of the world (Bruce et al., 2024). LLMs serve as the world simulator due to their generalizability across tasks. Specifically, The LLMs (i.e., GPT-3.5 and GPT-4) is evaluated to predict the state transitions, the game progress and scores with the given object, action, and score rules, where these rules are demonstrated to be crucial to the world model predictions (Wang et al., 2024). The world models with LLMs in (Xie et al., 2024) need to additionally identify the valid actions. We move a step further to ask the world model to propose the potential actions to complete the tasks (Observation 2). Both methods mainly focus on the prediction of the state, which may be not suitable for the evaluation of the world model for decision making (Observation 1).
World Models in LLMs.
The concept of world model also be explored in the deliberation reasoning of LLMs. Specifically, Reasoning via Planning (RAP) (Hao et al., 2023) leverage the planning methods (e.g., Monte Carlo Tree Search (MCTS)) with the world model with LLMs for plan generation and math reasoning, where LLMs need to predict the next state and the reward to guide the search. Tree of Thought (ToT) (Yao et al., 2023) implicitly leverage the LLMs as the world model to predict the next state and the reward for the search over different thoughts. Reason for future, act for now (RAFA) (Liu et al., 2023) combine the planning and reflection with the world model for complex reasoning tasks. However, these methods do not focus on the evaluation of the world models, and several interdependent modules are coupled with each other for completing the task (Observation 3).
3 Preliminaries
In this section, we will first introduce the preliminaries in decision making, the world model, search methods, and the recent practices that using LLM for decision making problems.
Markov Decision Process (MDP).
A decision making problem is usually represented as a Markov decision process (MDP) (Sutton & Barto, 2018), which is defined by the tuple $(S,A,T,R,\gamma)$, where $S$ is the state space, $A$ is the action space, $T:S\times A\rightarrow S$ is the transition dynamics, which specifies the next state $s^{\prime}$ given the current state $s$ and action $a$, $R:S\times A\rightarrow\mathbb{R}$ is the reward function, which specifies the agent’s reward given the current state $s$ and action $a$, and $\gamma$ is the discount factor. The agent’s policy is defined by $\pi_{\theta}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, parameterized by $\theta$, which takes the state $s$ as the input and outputs the action $a$ to be executed. The objective of the agent is to learn an optimal policy $\pi^{*}:=\arg\max_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|s_{0}\right]$ is the expected return and $s_{0}$ is the initial state.
Large Language Models (LLMs).
Large Language models (LLMs) learn from text data using unsupervised learning. LLMs optimize the joint probabilities of variable-length symbol sequences as the product of conditional probabilities by $P(x)=\prod\nolimits^{n}_{i=1}P(s_{i}|s_{1},...,s_{i-1})$, where $(s_{1},s_{2},...,s_{n})$ is the variable-length sequence of symbols. LLMs with billions of parameters have demonstrated impressive capabilities in various NLP tasks (Touvron et al., 2023; OpenAI, 2023).
World Models.
The world model $\Omega$ is introduced to predict the dynamics of the environment, thus supporting the decision making process. Specifically, the world model is trained or prompted to predict the next state $s^{\prime}$, the reward $r$, and the terminal function $d$, given the current state $s$ and action $a$. The world model can be a neural network specially trained on the environments (Hafner et al., 2019; Schrittwieser et al., 2020), which cannot generalize across different environments. Therefore, world models with LLMs emerge as a promising methods due to the generalizbility of LLMs, where the prompting (Xie et al., 2024), in-context learning (Wang et al., 2024), and even fine-tuning methods (Xiang et al., 2023; Lin et al., 2024) are used to transform the LLMs to the world models.
4 World Models with LLMs for Decision Making
Figure 3: The next state prediction and reward/terminal prediction are considered in (Wang et al., 2024). With the two predictions, we can complete policy verification. We introduce the action proposal of the world model and the policy planning can be completed with the policy verification and action proposal.
In this section, we introduce the world model with LLM for decision making. Specifically, we will introduce the next state prediction, the reward and terminal prediction. Then, we will introduce how the world model will be used to complete the considered three main tasks, i.e., policy verification, action proposal, and policy planning. We provide the relationship between the three main tasks and the two kinds of predictions in Figure 3 for better understanding of the rationale behind the three tasks.
The world model considered in this work mainly follows the design in (Wang et al., 2024), where the representation of the state includes the objects in the environments and their properties. The prompts to the LLM, e.g., GPT-4o, also include the object rules, the action rules, and the score rules, which provides the necessary knowledge of the environments for the LLM to make accurate predictions. For the next state prediction, we ask the LLM to predict the state changes, i.e., the change of the objects’ properties, which is demonstrated to be efficient for the prediction (Wang et al., 2024). With the predicted state changes, we can recover the full state for further predictions. For the reward/terminal prediction, the LLM needs to predict three features: i) gameScore: the reward received from the environment, ii) gameOver: whether the task is terminated, and iii) gameWon: whether the task is successfully completed or not. For the rules used for the prediction, we refer to (Wang et al., 2024) for more details and the code for generating the prompts is also provided in Appendix B.1.
Policy Verification.
Policy verification is one of the most straightforward way to apply the world model, which is motivated by Observation 1 to focus on the policy, rather than the predictions. The process for the policy verification is displayed in Algorithm 1. Specifically, given the environment env, the action sequence $\bm{a}$ with length $N$ to verify and the proportion of the action sequence to verify $\rho$, we will run the game for the first $\rho\cdot N$ steps (Line 1 of Algorithm 1), and leverage the world model to continue the last $(1-\rho)N$ steps (Line 1 of Algorithm 1). The returned results $r_{N},d_{N}$ will be compared with the true results from the environment to evaluate the performance of the world model for policy verification.
Action Proposal.
The action proposal is a novel task for world model, based on Observation 2, which is not considered in previous work. Therefore, we design the prompt for the world model to generate top-$K$ potential actions to complete the tasks. Specifically, we follow the representation of the state in the next state prediction, with the additional information: i) the examples of actions, and ii) the previous actions. The code to LLM for the action proposal is displayed in Appendix B.2. One key issue for the action proposal is that the action generated by the world model may not be valid for the game at the current state. Therefore, given the predicted action $a^{\prime}$ and the set of possible actions to be executed at the current state $A^{\prime}$, we leverage the text-embedding model (OpenAI, 2022) to query the most similar actions with the cosine similarity, i.e., $a^{*}=\arg\max\{\texttt{emb}(a^{\prime},a),\forall a\in A^{\prime}\}$.
1 Given the env, the action sequence $\bm{a}$ to verify with $N=\texttt{len}(\bm{a})$, $\rho$ the proportion of $\bm{a}$ to verify, the world model $\Omega$
[…]
3 for $t\in\{1,2,...,N-1\}$ do
4       if $t<\rho\cdot N$ then $s_{t+1},r_{t},d_{t}=\texttt{env}(a_{t})$;
5       else $s_{t+1},r_{t},d_{t}=\Omega(s_{t},a_{t})$;
6
return $r_{N},d_{N}$
Algorithm 1 Policy Verification
1 Given the env, the action sequence $\bm{a}$ with $N=\texttt{len}(\bm{a})$, $\rho$ the proportion of $\bm{a}$ for planning, the world model $\Omega$, the planning sequence $\bm{a^{\prime}}=[]$
2for $t\in\{1,2,...,\rho\cdot N\}$ do
3       $s_{t+1},r_{t}=\texttt{env}(a_{t}),\bm{a}^{\prime}.\texttt{append}(a_{t})$
4
5for $t\in\{\rho\cdot N,\cdots,(2-\rho)N\}$ do
6       $a_{t}=\Omega(s_{t}),s_{t+1},r_{t},d_{t}=\Omega(s_{t},a_{t}),\bm{a}^{\prime}.\texttt{append}(a_{t})$;
7       if $d_{t}$ then break;
8
9
return $\bm{a}^{\prime}$
Algorithm 2 Policy Planning
Policy Planning.
The policy planning task is motivated by Observation 3. The process of policy planning is displayed in Algorithm 2, which is built upon the policy verification and the action proposal (as depicted in Figure 3). Specifically, we execute the actions in the given $\bm{a}$ on the environment for $\rho N$ steps (Line 2 in Algorithm 2) and then plan for $2(1-\rho)N$ steps with the world model (Line 2 in Algorithm 2), where both the action to execute and the state transitions are generated by the world model. The returned action sequence $\bm{a}^{\prime}$ will be evaluated in the environment to verify the correctness. We note that only top-1 action is generated in Algorithm 2 for illustration, for the case where more actions are generated, we need to enumerate all possible outcomes or leverage advanced search methods, which will be tackled in future work.
5 Environments
Tasks.
We leverage the 31 diverse environments from (Wang et al., 2023) [^2] We note that there are 31 environments in (Wang et al., 2023), however, one environment, i.e., dish-washing, is used as the example for the world model, which is excluded for fair evaluation. with different tasks varying from daily tasks, e.g., washing clothes, to scientific tasks, e.g., forging keys, and different difficulty levels, i.e., steps to complete the tasks. The task suite is more diverse than the widely used environments, e.g., BabyAI (Chevalier-Boisvert et al., 2019) and MiniWob++ (Shi et al., 2017). A full list of the environments can be found in Appendix A.1.
Figure 4: Number of steps to complete the tasks
Rule-based Policies.
There are various randomness in the environments, including the specific tasks, e.g., the target color can be “orange”, “purple”, “green”, “black” in the paint task. However, for each task, only a single playthrough is provided in (Wang et al., 2024), which is not enough for a comprehensive evaluation of the world model for decision making. Therefore, we curate the rule-based policy for each environment and verify the correctness for 200 runs. The scripts for the rule-based policies are provided in Appendix A.2, which can help readers to understand the process to complete the tasks, as well as the complexities of tasks. We provide the statics of the number of steps to complete the tasks for 200 runs in Figure 4.333The names on the figure may differs from (Wang et al., 2023) for visualization, and please refer to Table 1 for correspondence.
6 Evaluations
In this section, we present the comprehensive evaluation of the world model for decision making over the diverse 31 environments on the three main tasks under various setttings. We use the advanced LLM models, i.e., GPT-4o and GPT-4o-mini, as the backbone LLM for the world model. We set the temperature of the LLM to be 0 to reduce the variance of the generation and all results are averaged over 30 runs.
6.1 Task I: Policy Verification
Figure 5: The accuracy of the world model to verify the correct policies
Evaluation Protocol.
Given the action sequence $\bm{a}$ generated by the rule-based policy, we leverage the world model to verify the last $\rho$ proportion of the policy, where $\rho\in\{0.25,0.5,0.75,1.0\}$. We note that when $\rho=1.0$, the world model will verify the full action sequence with only the initial observation of the environment. We say the verification of the policy is correct if all three features, i.e., gameScore, gameOver, and gameWon, are correct. We note that only correct policies are verified, as there are enormous wrong policies for an environment, which is useless for decision making. Furthermore, there would also be other action sequences to finish the tasks, where we cannot enumerate all policies to complete the tasks.
Evaluation Results.
The policy verification results are displayed in Figure 5. We observe that GPT-4o outperform GPT-4o-mini in most tasks and especially on the tasks which requires the domain knowledge, e.g., bandage, hang, and campfire. We also observe that with more steps of the verified policies, the performance gap between GPT-4o and GPT-4o-mini is increase. With larger proportion of the action sequences to verify, i.e., $\rho$ increase, the accuracy of the verification is decreased, which indicates that the accumulation of the errors in the world model, either on the next state prediction or the reward prediction, will influence the performance of the world model. This observation is consistent to the fact that the LLM may not perform well in long-term decision making tasks. We also observe that more steps to complete the tasks do not necessarily lead to the bad performance, which indicates that the domains of the tasks play a more important roles for the policy verification, i.e., for the tasks where the LLM has enough domain knowledge, the task would be easy even when the number of steps is large, e.g., conduct, stone, weigh and photo. We also provide the accuracy of the prediction of each feature in Appendix C, and we found that both GPT-4o and GPT-4o-mini performs wore for predicting the gameScore, while performs much better for predicting gameOver and gameWon, which indicates that the value prediction is more difficult for LLMs. During the experiments, we observe that both models usually output an empty dictionary of the verification result, which indicates the inabilities for the instruction following sometimes.
6.2 Task II: Action Proposal
Evaluation Protocol.
The action proposal requires the world model to generate the top-$K$ potential actions to complete the tasks, where $K\in\{1,2,3,5,10\}$. Specifically, given the action sequence $\bm{a}$ generated by the rule-based policies, we will let the world model to generate the potential actions with the states along with the path of $\bm{a}$ to complete the task. We say the action proposal is correct if the actions in $\bm{a}$ in the generated actions by the world model. The results of the accuracy are averaged over the steps over the action sequence and 30 runs of each environment. We also note that the action sequence $\bm{a}$ generated by the rule-based policy is not the only sequence to complete the task and we cannot enumerate all possible actions which can lead to the completion of the task. We note that the number of available actions in the environments is usually larger than 500, which brings difficulties to the traditional RL methods for training and indicate the necessity for the world model to generate the potential actions to facilitate the learning.
Figure 6: The accuracy of the world model to generate the potential actions
Evaluation Results.
The action proposal results are displayed in Figure 6. Overall, GPT-4o consistently outperforms GPT-4o-mini across different tasks and different values of $K$. With the number of steps to complete the tasks, GPT-4o maintains the better accuracy, while GPT-4o-mini shows a substantial drop of the accuracy. The performance gap between the two models is generally increased when the number of steps to complete the tasks increase. When $K=10$, the accuracy of the action proposal for GPT-4o is very high in most tasks, which indicates that the world model is capable to generate the relevant actions for completing the tasks while ignoring the irrelevant actions. Furthermore, we still observe that both models obtain lower values in the tasks requiring the domain knowledge, i.e., blood and conduct, which is consistent to the observation in (Wang et al., 2024) that LLMs (GPT-4) is more likely to make errors when scientific knowledge is needed. We also provide the step accuracy of the action proposal in Appendix D to illustrate the prediction of the relevant actions along with the steps. We observe that there are some key steps that has extremely low accuracies, which would be the most difficult steps for decision making.
6.3 Task III: Policy Planning
Evaluation Protocol.
The policy planning is based on the policy verification and the action proposal, as showed in Algorithm 2. Similar to the policy verification, we let $\rho\in\{0.25,0.5,0.75,1.0\}$ to vary the number of steps for the planning. We only consider the case with $K=1$, i.e., which generate the top-1 action with the given states predicted by the world model. Finally, we evaluate the planned policy $\bm{a}^{\prime}$ in the environment to verify the correctness. We note that when $K=1$, no advanced search method is needed, while when $K>1$, we cannot enumerate all possible outcomes for larger steps, e.g., 10. Besides, a critic is also needed to choose among the outcomes for verifying in the environments. Therefore, we only consider the case with $K=1$ and leave the case $K>1$ into future work.
Evaluation Results.
The policy planning results are displayed in Figure 7, where GPT-4o and GPT-4o-mini achieve comparable performance for the tasks with less steps and smaller values of $\rho$, e.g., $0.25$ and GPT-4o generally achieve better results in tasks with more steps. When the value of $\rho$ increases, the performance is generally decreasing. With the coupling of the policy verification and action proposal, we observe a more randomness of the performances of models over tasks and settings, which indicates that the necessity of decoupling the modules in decision making for comprehensive evaluation of the world model. During the experiments, we also observe the format errors of the outputs from both GPT-4o and GPT-4o-mini, which may interrupt the running of the experiments. Therefore, with the interaction of the different functionalities of the world model, the system is more unstable due to the unexpected in LLMs’ outputs.
Figure 7: The success rate of the world model to complete the tasks
7 Conclusions
World model is a key module for decision making and recent work leverage LLMs as the general world models. However, the evaluation of the world models with LLMs for decision making is far from satisfactory. In this work, we propose three main tasks, i.e., policy verification, action proposal, and policy planning, to evaluate the performance of the world model. We conduct the comprehensive evaluation of advanced LLMs, i.e., GPT-4o and GPT-4o-mini, on 31 diverse environments over the three main tasks under various settings. The key observations include: i) GPT-4o significantly outperforms GPT-4o-mini on the three main tasks, ii) the performance of the world model with LLM will be decreased for long-term decision-making tasks, and iii) the combination of different functionalities of the world model will brings unstability of the performance.
Appendix A Environments
A.1 Introduction of Tasks
Table 1: Environments (Wang et al., 2024)
Environments Task Description
mix-paint (paint) Your task is to use chemistry to create black paint.
blood-type (blood) Your task is to give a correct type of blood to the patient.
thermometer (thermo) Your task is to figure out the temperature of the water in the pot.
clean-energy (energy) Your task is to change all fossil-fuel power stations to use renewable energy while keeping the same capacity.
lit-lightbulb (bulb) Your task is to lit the light bulb.
scale-weigh (scale) Your task is to figure out the weight of the apple.
use-bandage (bandage) Your task is to put bandages on any cuts.
hang-painting (hang) Your task is to hang the picture of a girl (ID: 11) on the back wall (ID: 5).
multimeter (multi) Your task is to figure out the resistance of the resistor 0.
volume (volume) Your task is to figure out the volume of the green box.
sunburn (sunburn) It is a summer noon. The sky is clear. Your task is to take a ball from the beach and put it in the box in the house. Protect yourself from sunburn!
space-walk (space) Your task is to conduct a space walk.
sweep-floor (sweep) Your task is to clean the garbage on the ground to the garbage can.
volume-container (container) Your task is to figure out the volume of the glass.
bath-tub-water-temperature (bathtub) Your task is to make the temperature of the water in the bath tub to 35 - 40 Celsius degree by adding water from the taps. When you are done, take the action "bath".
conductivity (conduct) Your task is to figure out if the fork is conductive or not. If the fork is conductive, put it in the red box. Otherwise, put it in the black box.
make-campfire (campfire) Your task is to make a fire in the fire pit.
refrigerate-food (fridge) Your task is to prevent the foods from spoiling.
volume-stone (stone) Your task is to figure out the volume of the stone.
bird-life-cycle (bird) Your task is to hatch the egg and raise the baby bird.
balance-scale-weigh (weigh) Your task is to figure out the weight of the cube. Use the answer action to give your answer.
metal-detector (metal) Your task is to find the buried metal case on the beach. You win the game by putting the metal case in your inventory.
cooking (cooking) Your task is to prepare a meal following the instructions of the cook book.
make-ice-cubes (ice) Your task is to make ice cubes.
balance-scale-heaviest (heavy) Your task is to put all heaviest cubes into the box.
take-photo (photo) Your task is to take a nice picture of orange (ID: 4), using a camera with shutter speed of 1/2, aperture of 16, and iso of 1600.
plant-tree (plant) Your task is to plant the tree and water it.
boil-water (boil) Your task is to boil water.
forge-key (key) Your task is to forge a key to open the door.
inclined-plane (plane) Here are two inclined planes with the same angle. Your task is figure out which of the two inclined planes has the most friction. Focus on the inclined plane with the most friction after your experiment.
wash-clothes (washing) Your task is to wash the dirty clothes and dry them.
We note there are 32 environment in (Wang et al., 2023) and the dish-washing is selected as the example in the prompt, which is excluded for fair evaluation.
A.2 Code for Demo Actions Generation
Only one playthrough of the game is provided in (Wang et al., 2024), which is not enough due to the randomness in the environments. Therefore, we curate the
Code Sample 1: mix-paint
def get_demo_actions(self):
target_color = self.useful_info[0][0]
paint_names = self.useful_info[1]
color_dict = {
# "red": (1, 0, 0),
"orange": ["red", "yellow"],
# "yellow": (0, 1, 0),
"green": ["yellow", "blue"],
# "blue": (0, 0, 1),
"purple": ["red", "blue"],
"black": ["red", "yellow", "blue"],
paint_names_to_idx = {}
for paint_idx, paint_name in enumerate(paint_names):
paint_names_to_idx[paint_name] = paint_idx
color_mix_plan = color_dict[target_color]
demo_actions = []
to_idx = -1
for color_idx, color_mix in enumerate(color_mix_plan):
if color_idx == 0:
to_idx = paint_names_to_idx[color_mix]
continue
demo_actions.append(
"pour {} paint (ID: {}) in cup {} (ID: {})".format(
color_mix,
2 * (paint_names_to_idx[color_mix] + 1) + 1,
to_idx,
2 * (to_idx + 1),
demo_actions.append("mix cup {} (ID: {})".format(to_idx, 2 * (to_idx + 1)))
return demo_actions
Code Sample 2: blood-type
def get_demo_actions(self):
useful_info = self.useful_info[1]
return [
"give Type {} {} blood (ID: 3) to patient (ID: 2)".format(
useful_info[0], useful_info[1]
"take Type {} {} blood (ID: 3)".format(useful_info[0], useful_info[1]),
"give Type {} {} blood (ID: 3) to patient (ID: 2)".format(
useful_info[0], useful_info[1]
Code Sample 3: thermometer
def get_demo_actions(self):
demo_actions = [
"take thermometer (ID: 4)",
"use thermometer (ID: 4) on water (ID: 3)",
"answer {} Celsius degree".format(self.water_temperature),
return demo_actions
Code Sample 4: clean-energy
def get_demo_actions(self):
demo_actions = []
change_station = {
"sun": "solar farm",
"water": "hydroelectric power station",
"wind": "wind farm",
for region in self.regions:
demo_actions.append(
"change {} to {}".format(
region.name, change_station[region.properties["resource"]]
return demo_actions
Code Sample 5: lit-lightbulb
def get_demo_actions(self):
return [
"connect light bulb (ID: 2) terminal1 to red wire (ID: 3) terminal1",
"connect red wire (ID: 3) terminal2 to battery (ID: 6) anode",
"connect battery (ID: 6) cathode to black wire (ID: 4) terminal1",
"connect black wire (ID: 4) terminal2 to light bulb (ID: 2) terminal2",
Code Sample 6: scale-weigh
def get_demo_actions(self):
demo_actions = [
"take {}".format(self.useful_info[0].name),
"put {} on {}".format(self.useful_info[0].name, self.useful_info[1].name),
"look",
"answer {}g".format(self.target_weight),
return demo_actions
Code Sample 7: use-bandage
def get_demo_actions(self):
demo_actions = [
"open bandage box (ID: 8)",
"look",
"take bandage (ID: 9)",
return demo_actions + [
"put bandage (ID: 9) on {} (ID: 3)".format(self.useful_info[0])
Code Sample 8: hang-painting
def get_demo_actions(self):
demo_actions = [
"take nail (ID: 7)",
"take hammer (ID: 6)",
"hammer nail (ID: 7) on {} with hammer (ID: 6)".format(
self.target_wall.name
"take {}".format(self.target_picture.name),
"hang {} on nail (ID: 7)".format(self.target_picture.name),
return demo_actions
Code Sample 9: multimeter
def get_demo_actions(self):
demo_actions = [
"set multimeter (ID: 2) to resistance mode",
"connect multimeter (ID: 2) terminal1 to resistor {} (ID: 3) terminal1".format(
self.target_resistor_id
"connect multimeter (ID: 2) terminal2 to resistor {} (ID: 3) terminal2".format(
self.target_resistor_id
"look",
"answer {} ohm".format(self.target_resistance),
return demo_actions
Code Sample 10: volume
def get_demo_actions(self):
demo_actions = [
"take {}".format(self.useful_info[1].name),
"measure the length of the {} with the {}".format(
self.useful_info[0].name, self.useful_info[1].name
"measure the width of the {} with the {}".format(
self.useful_info[0].name, self.useful_info[1].name
"measure the height of the {} with the {}".format(
self.useful_info[0].name, self.useful_info[1].name
"answer {} cubic cm".format(self.target_box_volume),
return demo_actions
Code Sample 11: sunburn
def get_demo_actions(self):
return [
"use sunscreen (ID: 4)",
"move to beach (ID: 3)",
"look",
"take ball (ID: 8)",
"move to house (ID: 2)",
"put ball (ID: 8) in box (ID: 5)",
Code Sample 12: space-walk
def get_demo_actions(self):
return [
"put on space suit (ID: 7)",
"open inner door (ID: 5)",
"move to airlock (ID: 3)",
"look",
"close inner door (ID: 5)",
"open outer door (ID: 6)",
"move to outer space (ID: 4)",
Code Sample 13: sweep-floor
def get_demo_actions(self):
sweep_actions = []
for garbage in self.useful_info:
sweep_actions.append(
"sweep {} to dustpan (ID: 3) with broom (ID: 2)".format(garbage.name)
demo_actions = (
["take broom (ID: 2)", "take dustpan (ID: 3)"]
+ sweep_actions
"open garbage can (ID: 4)",
"dump dustpan (ID: 3) to garbage can (ID: 4)",
return demo_actions
Code Sample 14: volume-container
def get_demo_actions(self):
demo_actions = [
"take {}".format(self.useful_info[0].name),
"put {} in sink (ID: 2)".format(self.useful_info[0].name),
"turn on sink (ID: 2)",
"turn off sink (ID: 2)",
"take {}".format(self.useful_info[0].name),
"pour water in {} into {}".format(
self.useful_info[0].name, self.useful_info[1].name
"look",
"answer {} mL".format(self.target_water_container_volume),
return demo_actions
Code Sample 15: bath-tub-water-temperature
def get_demo_actions(self):
water_temp = self.useful_info[0]
cooling = [
"turn on cold tap (ID: 5)",
"turn off cold tap (ID: 5)",
"use thermometer (ID: 6) on water (ID: 3)",
hotting = [
"turn on hot tap (ID: 4)",
"turn off hot tap (ID: 4)",
"use thermometer (ID: 6) on water (ID: 3)",
if water_temp > 40:
cooling_times = (water_temp - 35) // 5
water_actions = cooling * cooling_times
elif water_temp < 35:
hotting_times = (40 - water_temp) // 5
water_actions = hotting * hotting_times
else:
water_actions = []
demo_actions = (
"take thermometer (ID: 6)",
"use thermometer (ID: 6) on water (ID: 3)",
+ water_actions
+ ["bath"]
return demo_actions
Code Sample 16: conductivity
def get_demo_actions(self):
demo_actions = [
"connect light bulb (ID: 2) terminal1 to red wire (ID: 3) terminal1",
"connect red wire (ID: 3) terminal2 to battery (ID: 6) anode",
"connect battery (ID: 6) cathode to black wire (ID: 4) terminal1",
"connect black wire (ID: 4) terminal2 to fork (ID: 7) terminal1",
"connect fork (ID: 7) terminal2 to blue wire (ID: 5) terminal1",
"connect blue wire (ID: 5) terminal2 to light bulb (ID: 2) terminal2",
"look",
"take fork (ID: 7)",
if self.useful_info[0]:
return demo_actions + ["put fork (ID: 7) in red box (ID: 8)"]
else:
return demo_actions + ["put fork (ID: 7) in black box (ID: 9)"]
Code Sample 17: make-campfire
def get_demo_actions(self):
return [
"take axe (ID: 4)",
"use axe (ID: 4) on tree (ID: 5)",
"look",
"use axe (ID: 4) on chopped down tree (ID: 5)",
"look",
"take firewood (ID: 5)",
"put firewood (ID: 5) in fire pit (ID: 2)",
"take match (ID: 3)",
"use match (ID: 3) on firewood (ID: 5)",
Code Sample 18: refrigerate-food
def get_demo_actions(self):
take_objects = []
put_objects = []
for food in self.useful_info:
take_objects.append("take {}".format(food.name))
put_objects.append("put {} in fridge (ID: 2)".format(food.name))
demo_actions = (
take_objects
+ ["open fridge (ID: 2)"]
+ put_objects
"close fridge (ID: 2)",
"look",
"look",
"look",
return demo_actions
Code Sample 19: volume-stone
def get_demo_actions(self):
demo_actions = [
"take measuring cup (ID: 4)",
"put measuring cup (ID: 4) in sink (ID: 2)",
"turn on sink (ID: 2)",
"turn off sink (ID: 2)",
"take measuring cup (ID: 4)",
"examine measuring cup (ID: 4)",
"take stone (ID: 3)",
"put stone (ID: 3) in measuring cup (ID: 4)",
"examine measuring cup (ID: 4)",
"answer {}".format(self.answer_volume),
return demo_actions
Code Sample 20: bird-life-cycle
def get_demo_actions(self):
return [
"sit on egg",
"sit on egg",
"sit on egg",
"sit on egg",
"sit on egg",
"feed young bird",
"feed young bird",
"feed young bird",
"feed young bird",
"feed young bird",
Code Sample 21: balance-scale-weigh
def get_demo_actions(self):
weight_list = [1, 1, 2, 5, 10]
weight_list_index = [False] * 5
def _find_combination():
remaining = self.cube_weight
# using 10
if remaining >= 10:
weight_list_index[-1] = True
remaining -= 10
# using 5
if remaining >= 5:
weight_list_index[-2] = True
remaining -= 5
# using 2
if remaining >= 2:
weight_list_index[-3] = True
remaining -= 2
if remaining > 0:
if remaining == 1:
weight_list_index[0] = True
if remaining == 2:
weight_list_index[0] = True
weight_list_index[1] = True
_find_combination()
weight_actions = []
for idx, weight_list_idx in enumerate(weight_list_index):
if weight_list_idx:
weight_actions += [
"take {}".format(self.useful_info[idx].name),
"put {} in right side of the balance scale (ID: 4)".format(
self.useful_info[idx].name
"look",
demo_actions = (
"take cube (ID: 10)",
"put cube (ID: 10) in left side of the balance scale (ID: 3)",
+ weight_actions
+ ["answer {}g".format(self.cube_weight)]
return demo_actions
Code Sample 22: metal-detector
def get_demo_actions(self):
agent_init_position = self.useful_info[0]
targe_position = self.useful_info[1]
direction = (
targe_position[0] - agent_init_position[0],
targe_position[1] - agent_init_position[1],
h_dir_list = (
["south"] * direction[0]
if direction[0] > 0
else ["north"] * (-direction[0])
v_dir_list = (
["east"] * direction[1] if direction[1] > 0 else ["west"] * (-direction[1])
dir_list = h_dir_list + v_dir_list
self.random.shuffle(dir_list)
detect_actions = [
"detect with metal detector (ID: 15)",
for dir_step in dir_list:
detect_actions.append("move {}".format(dir_step))
detect_actions.append("detect with metal detector (ID: 15)")
demo_actions = (
["take metal detector (ID: 15)", "take shovel (ID: 16)"]
+ detect_actions
+ ["look", "dig with shovel (ID: 16)", "look", "take metal case (ID: 11)"]
return demo_actions
Code Sample 23: cooking
def get_demo_actions(self):
cooking_actions = [
for cooking_item in self.receipt:
operations = self.receipt[cooking_item]
cooking_actions += ["take {}".format(cooking_item.name)]
for operation in operations:
if operation in ["slice", "dice", "chop"]:
cooking_actions += [
"{} {} with {}".format(
operation, cooking_item.name, self.useful_info["knife"].name
if operation in ["fry"]:
cooking_actions += [
"cook {} in {}".format(
cooking_item.name, self.useful_info["stove"].name
if operation in ["roast"]:
cooking_actions += [
"cook {} in {}".format(
cooking_item.name, self.useful_info["oven"].name
demo_actions = (
"take {}".format(self.useful_info["cook_book"].name),
"read {}".format(self.useful_info["cook_book"].name),
"take {}".format(self.useful_info["knife"].name),
+ cooking_actions
"prepare meal",
return demo_actions
Code Sample 24: make-ice-cubes
def get_demo_actions(self):
return [
"open freezer (ID: 2)",
"examine freezer (ID: 2)",
"take ice cube tray (ID: 3)",
"put ice cube tray (ID: 3) in sink (ID: 4)",
"turn on sink (ID: 4)",
"turn off sink (ID: 4)",
"take ice cube tray (ID: 3)",
"put ice cube tray (ID: 3) in freezer (ID: 2)",
"close freezer (ID: 2)",
"look",
"look",
"look",
Code Sample 25: balance-scale-heaviest
def get_demo_actions(self):
demo_actions = [
"take {}".format(self.useful_info[0][0].name),
"put {} in left side of the balance scale (ID: 3)".format(
self.useful_info[0][0].name
"take {}".format(self.useful_info[1][0].name),
"put {} in right side of the balance scale (ID: 4)".format(
self.useful_info[1][0].name
"look",
if len(self.useful_info) == 2:
if self.useful_info[0][1] > self.useful_info[1][1]:
# left is heavier
demo_actions.append("take {}".format(self.useful_info[0][0].name))
demo_actions.append(
"put {} in {}".format(
self.useful_info[0][0].name, self.answer_box.name
return demo_actions
elif self.useful_info[0][1] < self.useful_info[1][1]:
# right is heavier
demo_actions.append("take {}".format(self.useful_info[1][0].name))
demo_actions.append(
"put {} in {}".format(
self.useful_info[1][0].name, self.answer_box.name
return demo_actions
else:
demo_actions.append("take {}".format(self.useful_info[0][0].name))
demo_actions.append(
"put {} in {}".format(
self.useful_info[0][0].name, self.answer_box.name
demo_actions.append("take {}".format(self.useful_info[1][0].name))
demo_actions.append(
"put {} in {}".format(
self.useful_info[1][0].name, self.answer_box.name
return demo_actions
on_scale = [0, 1]
for i in range(2, len(self.useful_info)):
if self.useful_info[on_scale[0]][1] > self.useful_info[on_scale[1]][1]:
demo_actions += [
"take {}".format(self.useful_info[on_scale[1]][0].name),
"take {}".format(self.useful_info[i][0].name),
"put {} in right side of the balance scale (ID: 4)".format(
self.useful_info[i][0].name
"look",
on_scale[1] = i
else:
demo_actions += [
"take {}".format(self.useful_info[on_scale[0]][0].name),
"take {}".format(self.useful_info[i][0].name),
"put {} in left side of the balance scale (ID: 3)".format(
self.useful_info[i][0].name
"look",
on_scale[0] = i
max_weight = 0
if self.useful_info[on_scale[0]][1] > self.useful_info[on_scale[1]][1]:
demo_actions.append("take {}".format(self.useful_info[on_scale[0]][0].name))
demo_actions.append(
"put {} in {}".format(
self.useful_info[on_scale[0]][0].name, self.answer_box.name
max_weight = self.useful_info[on_scale[0]][1]
else:
demo_actions.append("take {}".format(self.useful_info[on_scale[1]][0].name))
demo_actions.append(
"put {} in {}".format(
self.useful_info[on_scale[1]][0].name, self.answer_box.name
max_weight = self.useful_info[on_scale[1]][1]
for cube, cube_mass in self.useful_info:
if cube_mass == max_weight:
demo_actions += [
"take {}".format(cube.name),
"put {} in {}".format(cube.name, self.answer_box.name),
return demo_actions
Code Sample 26: take-photo
def get_demo_actions(self):
speed = self.camera.properties["current_shutter_speed"]
iso = self.camera.properties["current_iso"]
aperture = self.camera.properties["current_aperture"]
target_aperture = self.useful_info[0]
target_speed = self.useful_info[1]
target_iso = self.useful_info[2]
adjust_actions = [
"focus {}".format(self.target_food.name),
if target_aperture > aperture:
adjust_actions += ["rotate aperture clockwise"] * (
target_aperture - aperture
elif target_aperture < aperture:
adjust_actions += ["rotate aperture anticlockwise"] * (
-target_aperture + aperture
else:
pass
if target_speed > speed:
adjust_actions += ["rotate shutter speed clockwise"] * (
target_speed - speed
elif target_speed < speed:
adjust_actions += ["rotate shutter speed anticlockwise"] * (
-target_speed + speed
else:
pass
if target_iso > iso:
adjust_actions += ["rotate iso clockwise"] * (target_iso - iso)
elif target_iso < iso:
adjust_actions += ["rotate iso anticlockwise"] * (-target_iso + iso)
else:
pass
demo_actions = ["take camera (ID: 2)"] + adjust_actions + ["press shutter"]
return demo_actions
Code Sample 27: plant-tree
def get_demo_actions(self):
return [
"take shovel (ID: 2)",
"dig with shovel (ID: 2)",
"take tree (ID: 8)",
"look",
"put tree (ID: 8) in hole (ID: 9)",
"inventory",
"put soil (ID: 10) in hole (ID: 9)",
"take jug (ID: 7)",
"put jug (ID: 7) in sink (ID: 5)",
"turn on sink (ID: 5)",
"turn off sink (ID: 5)",
"take jug (ID: 7)",
"pour water in jug (ID: 7) into soil (ID: 10)",
Code Sample 28: boil-water
def get_demo_actions(self):
return [
"take pot (ID: 4)",
"put pot (ID: 4) in sink (ID: 3)",
"examine sink (ID: 3)",
"turn on sink (ID: 3)",
"examine sink (ID: 3)",
"turn off sink (ID: 3)",
"take pot (ID: 4)",
"look",
"put pot (ID: 4) on stove (ID: 2)",
"examine stove (ID: 2)",
"turn on stove (ID: 2)",
"examine stove (ID: 2)",
"examine stove (ID: 2)",
"examine stove (ID: 2)",
Code Sample 29: forge-key
def get_demo_actions(self):
demo_actions = [
"take copper ingot (ID: 4)",
"put copper ingot (ID: 4) in foundry (ID: 3)",
"turn on foundry (ID: 3)",
"look",
"look",
"look",
"look",
"look",
"look",
"pour copper (liquid) (ID: 4) into key mold (ID: 6)",
"look",
"look",
"take copper key (ID: 4)",
"open door (ID: 5) with copper key (ID: 4)",
return demo_actions
Code Sample 30: inclined-plane
def get_demo_actions(self):
look_table = {
0.5: ["look"] * 5,
1: ["look"] * 5,
1.5: ["look"] * 5,
2: ["look"] * 5,
a1 = self.useful_info[0][0]
a2 = self.useful_info[0][1]
a1_look = look_table[a1]
a2_look = look_table[a2]
demo_actions = (
"take stopwatch (ID: 5)",
"take block (ID: 4)",
"put block (ID: 4) on inclined plane 1 (ID: 2)",
"activate stopwatch (ID: 5)",
+ a1_look
"deactivate stopwatch (ID: 5)",
"examine stopwatch (ID: 5)",
"reset stopwatch (ID: 5)",
"take block (ID: 4)",
"put block (ID: 4) on inclined plane 2 (ID: 3)",
"activate stopwatch (ID: 5)",
+ a2_look
"deactivate stopwatch (ID: 5)",
"examine stopwatch (ID: 5)",
if a1 > a2:
return demo_actions + ["focus on inclined plane 2 (ID: 3)"]
else:
return demo_actions + ["focus on inclined plane 1 (ID: 2)"]
Code Sample 31: wash-clothes
def get_demo_actions(self):
washing, drying, busketing = [], [], []
for cloth in self.dirty_clothes:
washing += [
"take {}".format(cloth.name),
"put {} in washing machine (ID: 2)".format(cloth.name),
drying += [
"take {}".format(cloth.name),
"put {} in dryer (ID: 3)".format(cloth.name),
busketing += [
"take {}".format(cloth.name),
"put {} in basket (ID: 12)".format(cloth.name),
for cloth in self.clean_clothes:
busketing += [
"take {}".format(cloth.name),
"put {} in basket (ID: 12)".format(cloth.name),
demo_actions = (
"open washing machine (ID: 2)",
+ washing
"use bottle of detergent (ID: 4) on washing machine (ID: 2)",
"close washing machine (ID: 2)",
"turn on washing machine (ID: 2)",
"wait",
"look",
"look",
"open washing machine (ID: 2)",
"open dryer (ID: 3)",
+ drying
"close dryer (ID: 3)",
"turn on dryer (ID: 3)",
"wait",
"look",
"look",
"open dryer (ID: 3)",
"take skirt (ID: 5)",
+ busketing
return demo_actions
A.3 Analysis of Demo Actions
Figure 8: Steps to complete the tasks
Appendix B Prompts for World Model
B.1 Prompt for Next State and Reward/Terminal Predictions
Code Sample 32: Code for Prompts of Generating Potential Actions.
prompt = (
"You are a simulator of a text game. Read the task description of a text game. "
"Given the current game state in JSON, "
[…]
prompt += (
"Your response should be in the JSON format. "
"It should have three keys: ’modified’, ’removed’, and ’score’. "
"The ’modified’ key stores a list of all the object states that are added or changed after taking the action. "
"Keep it an empty list if no object is added or modified. "
"The ’removed’ key stores a list of uuids of the objects that are removed. "
"Keep it an empty list if no object is removed. "
"The ’score’ key stores a JSON with three keys: "
"’score’, ’gameOver’, and ’gameWon’. "
"’score’ stores the current game score, "
"’gameOver’ stores a bool value on whether the game is over, "
[…]
last_action = "" if len(self.last_actions) == 0 else self.last_actions[-1]
max_UUID = importlib.import_module(self.game_name).UUID
if current_state is None:
current_state = get_state(self.game, last_action, max_UUID, self.game_name)
current_state_for_prompt = make_game_state(current_state)
max_uuid = current_state["max_UUID"]
else:
# print("use the predicted state")
current_state_for_prompt = current_state
max_uuid = len(current_state["game_state"])
# start adding examples
example_prompt = self.build_examples()
prompt += example_prompt
# end of adding examples
# Task
[…]
# load rules
obj_desc = preprocess_obj_desc(self.obj_rules[self.game_name])
action_desc = self.action_rules[self.game_name]
score_desc = self.score_rules[self.game_name]
[…]
prompt += obj_desc.strip()
[…]
prompt += action_desc.strip()
[…]
prompt += score_desc.strip()
[…]
# data_state, data_UUID_base, data_action = None, None, None
[…]
B.2 Prompts of Generating Potential Actions.
Code Sample 33: Code for Prompts of Generating Potential Actions.
prompt = (
"You are a simulator of a text game. "
"Read the task description and the descriptions of all game actions of a text game. "
"Given the current game state in JSON, and the previous actions that lead to the current game state, "
"you need to decide the most {} actions "
[…]
k
prompt += (
"Each of your action should in one phrase with one verb and the objects it operates on. "
[…]
prompt += (
"Your response should be in the JSON format. "
[…]
last_action = "" if len(self.last_actions) == 0 else self.last_actions[-1]
max_UUID = importlib.import_module(self.game_name).UUID
if current_state is None:
current_state = get_state(self.game, last_action, max_UUID, self.game_name)
current_state_for_prompt = make_game_state(current_state)
max_uuid = current_state["max_UUID"]
else:
# print("use the predicted state")
current_state_for_prompt = current_state
max_uuid = len(current_state["game_state"])
# start adding examples
# example_prompt = self.build_examples()
# prompt += example_prompt
# end of adding examples
# Task
[…]
# load rules
obj_desc = preprocess_obj_desc(self.obj_rules[self.game_name])
action_desc = self.action_rules[self.game_name]
score_desc = self.score_rules[self.game_name]
[…]
prompt += obj_desc.strip()
[…]
prompt += action_desc.strip()
[…]
prompt += score_desc.strip()
[…]
# data_state, data_UUID_base, data_action = None, None, None
[…]
if len(self.last_actions) == 0:
prompt += "There is no previous actions."
else:
[…]
"is" if len(self.last_actions) == 1 else "are"
for action in self.last_actions:
[…]
Appendix C Accuracy of Verifying Policies
Figure 9: The accuracy of the world model to verify the correct policies on score
Figure 10: The accuracy of the world model to verify the correct policies on gameOver
Figure 11: The accuracy of the world model to verify the correct policies on gameWon
Appendix D Step Accuracy of Action Proposal
Figure 12: Step correctness of the action proposal of GPT-4o-mini
Figure 13: Step correctness of the action proposal of GPT-4o